OpenStudy (anonymous):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

OpenStudy (anonymous):

Let x = first number y = second number You're given \[\begin{cases}5x+2y=8\\ 3x-7y=-22\end{cases}\] Do you know how to solve that?

OpenStudy (anonymous):

siths is 100% right

OpenStudy (anonymous):

yes solve it by making any one of the variable common in both the equations

OpenStudy (anonymous):

like multiply first eqn wid 7 and secnd eqn wid 2

OpenStudy (anonymous):

(5x+2y) x 7 = 8 x 7 35x + 14y = 56